OpenStudy (ledah):

Are 24, 45, and 51 the side measurements of a special right triangle?

OpenStudy (anonymous):

If they are, then Pythagoras's Theorem must apply; the square of the longest side is equal to the sum of the squares of the other two sides, in other words 51² = 24² + 45². But 51² = 2601. And 24² + 45² = 576 + 2025 = 2601, so yea they are the side measurements of a special right triangle.